OpenStudy (ejayy12):

What is the equation of the line that passes through (0, 2) and has a slope of 0? a.y=-2 b.y=2 c.x=2

OpenStudy (igreen):

Use point-slope form. \(\sf y - y_1 = m(x - x_1)\) (0, 2) x1 y1 m = 0 Plug them in and simplify

OpenStudy (tennis5518):

please put this in the math section and click on the openstudy feedback and choose the section you need help

OpenStudy (tennis5518):

:D

OpenStudy (igreen):

Can you do that? @ejayy12

rebeccaxhawaii (rebeccaxhawaii):

or\[slope=\frac{ y^2-y^1 }{ x^2-x^1 }\]

ILovePuppiesLol (ilovepuppieslol):

for future reference, put this in the math section http://openstudy.com/study#/groups/Mathematics

OpenStudy (ejayy12):

so is it A? y=-2

OpenStudy (igreen):

No, A is incorrect.

OpenStudy (igreen):

y - 2 = 0(x - 0) The right side is equal to 0, since we're distributing a 0. y - 2 = 0 Last step is to add 2 to both sides. y = 2